how many kg of hot water at 80C must me added to cold water at 15C in order to make a 1.5kg mixture at 25C? Neglect heat losses


1
Expert's answer
2022-05-16T17:03:05-0400

Heat balance says

"cm_1(t_3-t_1)=cm_2(t_2-t_3)"

"m_1+m_2=m"

"m_1(25-15)=m_2(80-25)\\
m_1+m_2=1.5"


Solution of this system of equations

"m_1=1.27\:{\rm kg}\\
m_2=0.23\:\rm kg"